OpenStudy (anonymous):

solve by completing the square. x^2+8x+____ a) 16 b) 4 c) 64 D)13 r^2-14r+____ a) -7 b) 49 c)-49 D)14

OpenStudy (dtan5457):

Same thing use the formula (b/2)^2

OpenStudy (dtan5457):

In x^2+8x+____ your b value is 8 r^2-14r+____ b value: -14

OpenStudy (dtan5457):

plug values to formula :)

OpenStudy (anonymous):

im not good at these things at all thanks soooo much :-*

OpenStudy (dtan5457):

what did you get as your answers? (if you got them now)

OpenStudy (anonymous):

64 for the first one

OpenStudy (dtan5457):

well.. (8/2)^2=(4)^2

OpenStudy (dtan5457):

is 4 times 4 64?

OpenStudy (anonymous):

no

OpenStudy (dtan5457):

then what is it? lol

OpenStudy (anonymous):

16 right

OpenStudy (dtan5457):

yes

OpenStudy (dtan5457):

try the 2nd one

OpenStudy (dtan5457):

(-7)^2=?

OpenStudy (anonymous):

49

OpenStudy (dtan5457):

good job :)

OpenStudy (anonymous):

thank you so much

OpenStudy (dtan5457):

np. feel free to medal as i do enjoy medals ;p

OpenStudy (anonymous):

answer is 16 because...we can write this as x^2+(2*4*x)+(16)=(x+4)^2

OpenStudy (anonymous):

how do i do medals

OpenStudy (dtan5457):

click "best response"

OpenStudy (dtan5457):

any moar questions??

OpenStudy (anonymous):

yeah ill write a new question after i close this one

OpenStudy (dtan5457):

ok. i'll go help you with that one too when you post it